# Open Biomedical Knowledge Graphs at Scale: Construction, Federation, and AI Agent Access with Samyama Graph Database

### 16 Mar 2026 v1

@@ Abstract
Biomedical knowledge is fragmented across siloed databases—Reactome for pathways, STRING for protein interactions, Gene Ontology for functional annotations, ClinicalTrials.gov for study registries, and dozens more. Researchers routinely download flat files from each source and write bespoke scripts to cross-reference them, a process that is slow, error-prone, and not reproducible. We present two open-source biomedical knowledge graphs—Pathways KG (118,686 nodes, 834,785 edges from 5 sources) and Clinical Trials KG (7,774,446 nodes, 26,973,997 edges from 5 sources)—built on Samyama, a high-performance graph database written in Rust.
Our contributions are threefold. First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch Cypher loading, and portable snapshot export. Second, we demonstrate cross-KG federation: loading both snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “Which biological pathways are disrupted by drugs currently in Phase 3 trials for breast cancer?”—a query that neither KG can answer alone. Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries without manual tool authoring.
All data sources are open-license (CC BY 4.0, CC0, OBO). Snapshots, ETL code, and MCP configurations are publicly available. The combined federated graph (7.89M nodes, 27.8M edges) loads in 76 seconds on commodity hardware (Mac Mini M4, 16GB RAM), and the signature cross-KG query—“which pathways are disrupted by drugs in Phase 3 breast cancer trials?”—returns validated results in 2.1 seconds.
Keywords: Knowledge Graphs, Biomedical Data Integration, Graph Databases, Cross-KG Federation, Model Context Protocol, Clinical Trials, Biological Pathways, OpenCypher.
@@ 1 Introduction
However, constructing biomedical KGs at scale remains challenging. Prior efforts such as Bio2RDF (Callahan et al., 2013), Hetionet (Himmelstein et al., 2017), and the Clinical Knowledge Graph (Santos et al., 2022) have made significant progress, but each faces limitations: Bio2RDF requires a dedicated SPARQL endpoint infrastructure, Hetionet is a static dataset without an update pipeline, and the Clinical Knowledge Graph targets a single data source.
@@
    Reproducible KG construction. We present ETL pipelines for two large-scale biomedical KGs—Pathways KG (5 sources, 119K nodes) and Clinical Trials KG (5 sources, 7.7M nodes)—built on Samyama Graph Database using a common pattern: download, parse, deduplicate, batch-load via Cypher, and export as portable .sgsnap snapshots.
@@
    Cross-KG federation. We show that loading multiple snapshots into a single graph tenant enables property-based federation—joining on shared identifiers (UniProt accessions, DrugBank IDs, NCBI Gene IDs) without ETL-time merging. We demonstrate five cross-KG query patterns spanning translational medicine.
@@
    Schema-driven AI agent access. Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration that auto-generates typed tools from the graph schema. LLM agents can query the KG via natural language without hand-coded tool definitions.
All code, snapshots, and MCP configurations are open-source [^1] Pathways KG: https://github.com/samyama-ai/pathways-kg; Clinical Trials KG: https://github.com/samyama-ai/clinicaltrials-kg; Snapshots: https://github.com/samyama-ai/samyama-graph/releases.
@@ 2.1 Biomedical Knowledge Graphs
Bio2RDF (Callahan et al., 2013) converts life science databases to RDF and serves them via SPARQL endpoints. The RDF approach offers semantic interoperability but introduces complexity (OWL reasoning, blank nodes) and performance overhead for traversal-heavy queries.
@@ 3 Data Sources
Table 1: Data sources for the two biomedical KGs. All are open-license and human-specific (organism 9606 where applicable).
@@ 5.1 Motivation
The Pathways KG knows molecular biology—which proteins interact, what pathways they participate in. The Clinical Trials KG knows translational medicine—which drugs are in trials, what conditions they treat. Neither alone can answer:
“Which biological pathways are disrupted by drugs currently in Phase 3 trials for breast cancer?”
@@ 5.2 Join Points
Three entity types appear in both KGs with matching identifiers:
Table 4: Cross-KG join points.
Entity Pathways Property Clinical Trials Property Identifier
Protein Protein.uniprot_id Protein.uniprot_id UniProt accession
Drug Drug.drugbank_id Drug.drugbank_id DrugBank ID
Gene Gene.gene_id Gene.gene_id NCBI Gene ID
@@ 7 Evaluation
We evaluate construction time, query performance, and federation correctness on a Mac Mini M4 (Apple M4, 10 cores, 16GB RAM).
@@ 7.1 Construction Performance
Table 7: KG construction performance.
KG Nodes Edges Snapshot Size Import Time
Pathways KG 118,686 834,785 9 MB 1 s
Clinical Trials KG 7,774,446 26,973,997 711 MB 72 s
Drug enrichment (32 drugs) +33 +235 3 s
Combined 7,893,165 27,809,017 720 MB 76 s
Snapshot import uses gzip-compressed JSON-lines format. The Pathways KG loads in 1 second; the Clinical Trials KG, with its 7.8M nodes, loads in 72 seconds. A targeted drug enrichment step adds 32 breast cancer drug nodes with 36 TARGETS edges and 199 CODED_AS_DRUG edges in 3 seconds. The full pipeline—from empty tenant to federated graph—completes in 76 seconds.
@@ 7.2 Query Performance
Table 8: Query latency on Mac Mini M4.
@@
Label distribution (15 labels) Combined 15 rows 7.8 s
Proteins with uniprot_id Combined 1 row 0.5 s
Top pathways by protein count Pathways 10 rows 85 ms
PPI hub proteins (TP53: 739) Pathways 5 rows 1.4 s
Breast cancer trial count Clinical 5 rows 1.2 s
Drug $\rightarrow$ Pathway (6-hop) Combined 10 rows 2.1 s
The signature federated query—“pathways disrupted by breast cancer trial drugs”—traverses 6 hops across both KGs and returns in 2.1 seconds. Label distribution over the full 7.9M-node graph takes 7.8 seconds (full scan). Individual KG queries execute in under 1.5 seconds.
@@ 10 Conclusion
We presented two open-source biomedical knowledge graphs—Pathways KG and Clinical Trials KG—built on Samyama Graph Database. Together they contain 7.89 million nodes and 27.8 million edges from 10 public data sources. We demonstrated that loading both snapshots into a single graph tenant, combined with a targeted drug-target enrichment step (32 drugs, 36 TARGETS edges), enables cross-KG federated queries that bridge molecular biology and translational medicine. The 6-hop query “pathways disrupted by Phase 3 breast cancer drugs” returns biologically validated results—Signal Transduction, Immune System, Cell Cycle—in 2.1 seconds on commodity hardware. We introduced schema-driven MCP server generation for LLM agent access to graph queries.
All code, snapshots, ETL pipelines, and MCP configurations are open-source under Apache License 2.0. Researchers can reproduce the federation—from snapshot import to cross-KG queries—in under 2 minutes on commodity hardware.

### 17 Mar 2026 v2

@@ Abstract
Biomedical knowledge is fragmented across siloed databases—Reactome for pathways, STRING for protein interactions, Gene Ontology for functional annotations, ClinicalTrials.gov for study registries, DrugBank for drug vocabularies, DGIdb for drug–gene interactions, SIDER for side effects, and dozens more. Researchers routinely download flat files from each source and write bespoke scripts to cross-reference them, a process that is slow, error-prone, and not reproducible. We present three open-source biomedical knowledge graphs—Pathways KG (118,686 nodes, 834,785 edges from 5 sources), Clinical Trials KG (7,774,446 nodes, 26,973,997 edges from 5 sources), and Drug Interactions KG (32,726 nodes, 191,970 edges from 3 sources)—built on Samyama, a high-performance graph database written in Rust.
Our contributions are threefold. First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch loading (both Python Cypher and Rust native loaders), and portable snapshot export. Second, we demonstrate cross-KG federation: loading all three snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”—a query that no single KG can answer alone. Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries. We evaluate domain-specific MCP tools against text-to-Cypher and standalone GPT-4o on a new BiomedQA benchmark (40 pharmacology questions), achieving 98% accuracy vs. 0% for text-to-Cypher and 75% for standalone GPT-4o.
All data sources are open-license (CC BY 4.0, CC0, OBO, public domain). Snapshots, ETL code, and MCP configurations are publicly available. The combined federated graph (7.9M nodes, 28M edges) loads in approximately 3 minutes from portable snapshots on commodity cloud hardware (AWS g4dn.4xlarge, 62 GB RAM), and cross-KG queries complete in 80 ms–4 s.
Keywords: Knowledge Graphs, Biomedical Data Integration, Graph Databases, Cross-KG Federation, Model Context Protocol, Clinical Trials, Biological Pathways, Drug Interactions, Pharmacogenomics, OpenCypher.
@@ 1 Introduction
However, constructing biomedical KGs at scale remains challenging. Prior efforts such as Bio2RDF (Belleau et al., 2008), Hetionet (Himmelstein et al., 2017), and the Clinical Knowledge Graph (Santos et al., 2022) have made significant progress, but each faces limitations: Bio2RDF requires a dedicated SPARQL endpoint infrastructure, Hetionet is a static dataset without an update pipeline, and the Clinical Knowledge Graph targets a single data source.
@@
    Reproducible KG construction. We present ETL pipelines for three biomedical KGs—Pathways KG (5 sources, 119K nodes), Clinical Trials KG (5 sources, 7.7M nodes), and Drug Interactions KG (3 sources, 33K nodes)—built on Samyama Graph Database using a common pattern: download, parse, deduplicate, load (Python Cypher or Rust native), and export as portable .sgsnap snapshots. The Rust native loader creates the Drug Interactions KG (32K nodes, 192K edges) in under 1 second.
@@
    Cross-KG federation. We show that loading multiple snapshots into a single graph tenant enables property-based federation—joining on shared identifiers (UniProt accessions, DrugBank IDs, gene names) without ETL-time merging. We demonstrate cross-KG query patterns spanning molecular biology, translational medicine, and pharmacogenomics.
@@
    Schema-driven AI agent access. Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration with domain-specific tools. We evaluate MCP tools against text-to-Cypher (GPT-4o) and standalone GPT-4o on a new BiomedQA benchmark: MCP tools achieve 98% accuracy (39/40) vs. 0% for text-to-Cypher and 75% for standalone GPT-4o.
All code, snapshots, and MCP configurations are open-source [^1] Pathways KG: https://github.com/samyama-ai/pathways-kg; Clinical Trials KG: https://github.com/samyama-ai/clinicaltrials-kg; Drug Interactions KG: https://github.com/samyama-ai/druginteractions-kg; Snapshots: https://github.com/samyama-ai/samyama-graph/releases.
@@ 2.1 Biomedical Knowledge Graphs
Bio2RDF (Belleau et al., 2008) converts life science databases to RDF and serves them via SPARQL endpoints. The RDF approach offers semantic interoperability but introduces complexity (OWL reasoning, blank nodes) and performance overhead for traversal-heavy queries.
@@ 3 Data Sources
Table 1: Data sources for the three biomedical KGs. All are open-license and human-specific (organism 9606 where applicable).
@@
Drug Int. DrugBank CC0 Drug vocabulary (19K drugs) CC0 3 MB
DGIdb Drug–gene interactions (38K edges) Open 24 MB
SIDER Side effects & indications $C$-BY-SA 22 MB
@@ 4.3 Clinical Trials KG
4.4 Drug Interactions KG
The Drug Interactions KG models drug–gene interactions, side effects, and indications with 4 node labels and 3 edge types (Table 4).
Table 4: Drug Interactions KG schema.
Node Label Key Property Count
Drug drugbank_id 19,842
Gene gene_name 4,182
SideEffect meddra_id 5,858
Indication meddra_id 2,844
Total nodes 32,726
Edge Type Count
INTERACTS_WITH_GENE 38,033
HAS_SIDE_EFFECT 139,193
HAS_INDICATION 14,744
Total edges 191,970
The ETL consists of two phases. Phase 1 (DrugBank + DGIdb): 19,842 Drug nodes are created from DrugBank CC0 vocabulary CSV, with all synonyms indexed for cross-source name matching (52,154 synonyms). DGIdb interactions yield 4,182 Gene nodes and 38,033 INTERACTS_WITH_GENE edges with interaction types (inhibitor, activator, etc.). Phase 2 (SIDER): STITCH compound IDs are mapped to DrugBank names via synonym lookup, yielding 5,858 SideEffect nodes with 139,193 HAS_SIDE_EFFECT edges and 2,844 Indication nodes with 14,744 HAS_INDICATION edges.
A Rust native loader (druginteractions_loader.rs) performs the full ETL in 928 ms using direct GraphStore API calls (no Cypher parsing), compared to $\sim$}50 minutes via the Python HTTP loader.
@@ 5.1 Motivation
The Pathways KG knows molecular biology—which proteins interact, what pathways they participate in. The Clinical Trials KG knows translational medicine—which drugs are in trials, what conditions they treat. The Drug Interactions KG knows pharmacology—which genes a drug targets, what side effects it causes. No single KG can answer:
“For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”
@@ 5.2 Join Points
Shared entity types with matching identifiers enable cross-KG joins:
Table 5: Cross-KG join points across the three biomedical KGs.
Entity From KG To KG Join Property Identifier
Gene/Protein Drug Int. Pathways gene_name = name Gene symbol
Drug Drug Int. Clin. Trials name = Intervention.name Drug name
Drug Drug Int. Clin. Trials drugbank_id DrugBank ID
Protein Clin. Trials Pathways uniprot_id UniProt accession
@@ 7 Evaluation
We evaluate construction time, query performance, and AI agent access on an AWS g4dn.4xlarge instance (16 vCPU AMD EPYC, 62 GB RAM, NVIDIA A10G GPU).
@@ 7.1 Construction Performance
Table 8: KG construction performance (snapshot import on AWS g4dn.4xlarge).
KG Nodes Edges Snapshot Import Rust ETL
Pathways KG 118,686 834,785 9 MB 3.4 s
Drug Interactions KG 32,726 191,970 1.9 MB 0.7 s 0.9 s
Clinical Trials KG 7,774,446 26,973,997 711 MB 177 s
Combined 7,925,858 28,000,752 722 MB 181 s
Snapshot import uses gzip-compressed JSON-lines format. The Pathways KG loads in 3.6 seconds; the Drug Interactions KG in 0.7 seconds; and the Clinical Trials KG, with its 7.8M nodes, loads in 174 seconds. The Drug Interactions KG also has a Rust native loader that constructs the full graph from source files (DrugBank CSV, DGIdb TSV, SIDER TSV) in 928 ms—orders of magnitude faster than the Python HTTP-based ETL. The combined federated graph (7.9M nodes, 28M edges) uses 33 GB RAM (53% of 62 GB available on the AWS instance).
@@ 7.2 Query Performance
Table 9: Query latency on AWS g4dn.4xlarge (7.9M nodes loaded).
@@
Drug gene targets (Metformin) Drug Int. 5 rows 83 ms
Shared gene targets (2 drugs) Drug Int. 1 row 84 ms
Top pathways by protein count Pathways 10 rows 97 ms
PPI partners (TP53: Q9UQ61) Pathways 10 rows 96 ms
Side effects of Warfarin Drug Int. 10 rows 82 ms
Drugs for diabetes $\rightarrow$ gene targets Drug+Path 10 rows 3.6 s
EGFR drugs $\rightarrow$ side effects Drug Int. 10 rows 3.7 s
Cross-KG: PTGS1 drugs + side effects All 3 10 rows 3.7 s
Simple single-KG queries (drug lookups, gene targets, side effects) complete in 80–100 ms. Multi-hop cross-KG joins (e.g., “drugs targeting PTGS1 and their most common side effects”) complete in 3–4 s on the full 7.9M-node graph.
7.3 BiomedQA Benchmark
We introduce BiomedQA, a benchmark of 40 pharmacology questions across 7 categories over the three federated KGs. We compare three approaches: domain-specific MCP tools (parameterized Cypher templates), text-to-Cypher (GPT-4o generates Cypher from a schema prompt with 3 examples), and standalone GPT-4o (no database access).
Table 10: BiomedQA results (40 questions, 7.9M nodes, 3 federated KGs).
Approach Accuracy Avg Latency Avg Tokens
GPT-4o standalone 30/40 (75%) 2,474 ms 213
Text-to-Cypher (GPT-4o) 0/40 (0%) 986 ms 548
MCP tools 39/40 (98%) 651 ms 0
The text-to-Cypher result (0%) is striking: GPT-4o generates syntactically plausible Cypher that silently returns empty results due to schema mismatches (hallucinated property names, incorrect multi-MATCH patterns). MCP tools eliminate this class of failure entirely. The benchmark and evaluation scripts are open-source.
@@ 10 Conclusion
We presented three open-source biomedical knowledge graphs—Pathways KG (119K nodes from 5 sources), Clinical Trials KG (7.7M nodes from 5 sources), and Drug Interactions KG (33K nodes from 3 sources)—built on Samyama Graph Database. Together they contain 7.9 million nodes and 28 million edges from 13 public data sources. Loading all three snapshots into a single graph tenant enables cross-KG federated queries that bridge molecular biology, translational medicine, and pharmacogenomics. Cross-KG queries complete in 80 ms–4 s on commodity hardware (62 GB RAM). A Rust native loader constructs the Drug Interactions KG in under 1 second.
We introduced domain-specific MCP tools for LLM agent access and evaluated them on a new BiomedQA benchmark (40 pharmacology questions): MCP tools achieve 98% accuracy vs. 0% for text-to-Cypher and 75% for standalone GPT-4o, demonstrating that for domain-specific data access, the LLM’s role should be tool selection and argument extraction, not query generation.
All code, snapshots, ETL pipelines, benchmarks, and MCP configurations are open-source. Researchers can reproduce the full 3-KG federation—from snapshot import to cross-KG queries—in under 3 minutes.